# WaterVG: Waterway Visual Grounding based on Text-Guided Vision and mmWave Radar

### v1

@@ 3.2. Dataset Statistics
Figure 4 presents the statistics of WaterVG from various perspectives. Figure 4(a) shows the waterway scenario proportion in WaterVG, including canal, river, lake and moat. Figure 4(b) presents the proportion of sentence patterns for prompts, including statement (egocentric), question, statement (imperative) and exclamation. Figure 4(c) is the prompt type proportion, containing 4 main types of query by exact features, query by partial features, query by number and query by category. Figure 4(d) illustrates the distribution of text prompt lengths, ranging from 2 to 40, encompassing both simple phrases and detailed long sentences. The word cloud in Figure 4(e) of WaterVG displays high-frequency words, including categories, appearance, location, distance, motion, spatial relationships, etc., reflecting the richness and comprehensiveness of the dataset text prompt. Figure 4(f) shows the co-occurrence matrix of referred targets for various categories, which partly reflects the category diversity of the targets described in each of our samples. Figure 4(g) illustrates the distribution statistics of referred target number per sample. In WaterVG, each text prompt refers to 1-15 targets, with the majority consisting of 1-2 targets. Figure 4(h) presents the distribution of referred target categories, including piers, buoys, sailors, ships, boats, pleasure boats and kayaks, a total of 34,950 referred targets. The above demonstrates the diversity and comprehensiveness of our WaterVG dataset in terms of visual grounding in waterway navigation. This is primarily reflected in the diversity of waterway scenes, richness of textual prompts in terms of content, patterns, and types, diversity of mentioned target categories, and flexibility in the number of mentioned targets.
@@ 4. Methods
We propose a one-stage multi-task visual grounding network called Potamoi to simultaneously conduct two grounding tasks, REC and RES. We first introduce the overall pipeline of Potamoi as Figure 5 shows, and then illustrate our proposed Phased Heterogeneous Modality Fusion (PHMF), including (a) Adaptive Radar Weighting (ARW) and (b) Multi-Head Slim Cross Attention (MHSCA) in detail.
@@ 4.1. Overall Pipeline
Given an input triplet of an image ($I\in R^{3\times I_{H}\times I_{W}}$), a frame of radar point clouds ($R$) and a text prompt $T$, feature encoders for three modalities are for mapping them to high-dimension latent space. Firstly, the image encoder adopts FastViT-SA12 (Vasu et al., 2023) with the cascade of reparametrized RepMixer and Multi-Head Self-Attention (MHSA). Secondly, for 3D radar point clouds, we primarily project point clouds from radar coordinate to image coordinate, and obtain a three-channel radar point cloud feature map called RVP maps. Each point cloud contains the target Range (R), compensated Velocity (V) and reflected Power (P). For the RVP map with sparse point cloud features, we adopt the combination of convolution, batch normalization and ReLU activation as the basic block. Each stage of radar backbone contains three basic blocks, where the convolution kernels are $3\times 3$, $1\times 1$ and $3\times 3$ with the stride of $1$, $1$ and $2$ for local feature extraction, feedforward and downsampling. Whatever image or radar encoder, there are four stages and one stem layer, and the stem layer downsamples the spatial size four times. The feature dimension of stage $i\in\{1,2,3,4\}$ is $(64\times 2^{i-1})\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$.
Importantly, the last three stage features from the radar encoder go through three ARW modules and fuse with the same-dimensional feature maps from the image encoder by element-wise addition in the vision-radar fusion network, whose outputs are three-stage sensor features. These sensor features then act as queries to fuse with the keys and values of textual features from the pretrained ALBERT (Lan et al., 2020) by our proposed MHSCA and output three-stage fusion features. Then, a Path Aggregation Network (PANet) (Li et al., 2022) is to fuse features with various receptive fields, and outputs four-stage pyramid features from $P2$ to $P5$.
@@
We feed all four pyramid features to REC heads, which are transplanted and adapted from YOLOv8 decouple head (Jocher et al., 2023). Each pyramid feature corresponds to one REC head, containing a confidence branch and a bounding box regression branch. The shape of REC output for the stage $i\in\{2,3,4,5\}$ is $(1+4\times RegMax)\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$, where $1$ is the confidence score and $4$ is four coordinates of bounding box. $RegMax$ is the hyperparameter of Distributed Focal Loss (DFL) (Li et al., 2020). Furthermore, following the paradigm of PP-LiteSeg (Peng et al., 2022), we feed $P3$, $P4$ and $P5$ upon SPPM to RES head, which is based on Flexible Lightweight Decoder (FLD) (Peng et al., 2022). The output shape of RES head is $2\times I_{H}\times I_{W}$, including referred foreground and unreferred background categories.
@@ 4.2. Phased Heterogeneous Modality Fusion
We adopt a divide-and-conquer strategy for three modalities, including Adaptive Radar Weighting and Multi-Head Slim Cross Attention, which are vision-radar and text-sensor fusion in essence.
@@
(a) Adaptive Radar Weighting. Firstly, radar, as an auxiliary modality, can provide target prior information on distance, speed (motion), and target reflectivity to assist in distinguishing target categories. When combined with images for text queries, two considerations arise regarding the RVP map. Firstly, not all text prompts contain distance or motion information, corresponding to the channel features of the RVP map. Secondly, radar can be affected by clutter and multi-path effects, leading to generation of useless clutter point clouds in environments, corresponding to the spatial features of the RVP map. To adaptively balance the role of RVP map in visual grounding, we propose a module called Adaptive Radar Weighting (ARW) to adaptively weigh the RVP map. As Figure 6 presents, assuming the radar feature is $F_{R}\in R^{C\times H\times W}$, we exert a linear projection with GeLU activation on it to obtain $\hat{F_{R}}\in R^{C\times H\times W}$. The resulting feature map is then prepared for adaptive channel attention matrix $W_{CA}\in R^{C}$ by applying adaptive average pooling to collapse spatial dimensions and computing attention weights by softmax. $\hat{F_{R}}$ is position-wise multiplication with $W_{CA}$ to obtain the channel-weighting feature $F_{R}^{CW}\in R^{C\times H\times W}$. Then spatial weighting matrix $W_{SA}\in R^{1\times H\times W}$ is achieved by convolving $F_{R}^{CW}$ with a GELU activation and a 1 $\times$ 1 convolution, followed by element-wise multiplication with the original features and obtain the final radar feature $F_{R}^{ARW}\in R^{C\times H\times W}$ with both channel and spatial adaptive weighting. Equation 1 and 2 present the whole process above.
@@
Figure 7. The structure of Slim Cross Attention.
(b) Multi-Head Slim Cross Attention. After obtaining three-stage sensor features ($F^{2}_{S}$, $F^{3}_{S}$ and $F^{4}_{S}$) containing the contextual features of vision and prior target features captured by radar, there are three linear projection branches from ALBERT to align the channel dimensions of three-stage sensor features. To effectively fuse two heterogeneous modalities with a relatively low cost, we propose Multi-Head Slim Cross Attention (MHSCA) with linear complexity. Given a sensor feature $F_{S}\in R^{C\times H\times W}$ and a text feature $F_{T}\in R^{L\times C}$, we first exert learnable encodings (Dosovitskiy et al., 2020)(Vaswani et al., 2017), $P_{I}$ and $P_{T}$ on them, respectively. Then we adopt 2D depth-wise convolution with $1\times 1$ kernel ($W_{Q}$) to obtain the sensor query ($Q$) while two 1D depth-wise convolutions ($W_{K}$ and $W_{V}$) are for generate key ($K$) and value ($V$). Equation 3 presents the aforementioned process.
$Q=F_{S}W_{Q}+P_{S},K=F_{T}W_{K}+P_{T},V=F_{V}W_{V}+P_{T},$
(3) $Q\in R^{C\times H\times W},K\in R^{L\times C},V\in R^{L\times C},$
Further, we employ adaptive pooling to adaptively downsample the sensor query $Q$ (Han et al., 2023), preserving its essential features while reducing its spatial size to obtain the sensor agent $A\in R^{C\times h\times w}$. $A$, as the agent of $Q$, is firstly flatten along spatial dimension as $A\in R^{n\times C}$ and conducts softmax attention ($\sigma(\cdot)$) with $K$ and $V$ to calculate the agent-conditional Value matrix $V_{A}$. (Equation 4)
$A=Flat(A)^{T},A\in R^{n(h\times w)\times C},$
(4) $V_{A}=\sigma(A,K,V)=\frac{AK^{T}}{\sqrt{d}}\cdot V,$
@@
Then, $V_{A}$ is forwarded to the second softmax attention, broadcasting the inter-modal similarity embedded in text feature to the sensor query $Q$, where $A$ acts as the key and $V_{A}$ is the value (Equation 5). Here $Q$ is also firstly flattened as $Q\in R^{N\times C}$, where $N>n$.
$Q=Flat(Q)^{T},Q\in R^{N(H\times W)\times C},$
(5) $\sigma(Q,A,V_{A})=\frac{QA^{T}}{\sqrt{d}}\cdot V_{A},$
@@
(6) $F_{TS}=(Reshape(\sigma(Q,A,V_{A}))+F_{S})W,F_{TS}\in R^{C\times H\times W},$
@@ 4.3. Training Objectives
We reformulate the REC and RES tasks as the bounding box regression and semantic segmentation with one category. For both two tasks, the area occupied by the referred objects in the image is much smaller than the background area. Equation 7 presents the REC loss. We introduce Distributed Focal Loss (DFL) (Li et al., 2020) to alleviate the unbalanced distribution between two areas. Meanwhile, CIoU (Zheng et al., 2020) is added to refinedly regress the bounding box size by taking its aspect ratio into account.
@@
For RES task, we combine the dice loss with focal loss to jointly subtly optimize the mask prediction.
@@
We empirically argue that in the joint optimization process of REC and RES, the loss functions are not in the same order of magnitude. To avoid the domination of optimization by the task with a large loss, we introduce uncertainty weighting (Kendall et al., 2018) to balance multi-task optimization. Equation 9 demonstrates the loss function during the joint optimization of REC and RES.
@@
Figure 8. Images with projected radar point clouds (PCs) (first row) and prediction results by Potamoi (second row) and MCN (Luo et al., 2020) (third row) under different situations. Red bounding boxes are REC predictions by text prompt while translucent red masks are RES predictions. Blue PCs denote clutter while other represent targets. The small gray letter on the PC refers to the distance (m) from targets to the ego USV.
@@ 5.1. Settings of Dataset and Implementation
Datasets. Our proposed WaterVG is for training (9,254), validation (578) and testing (1,736) on all models. Besides, our WaterVG includes two benchmarks, Fusion-based Visual Grounding (FVG) and Vision-based Visual Grounding (VVG).
Models. For the REC, MattNet (Yu et al., 2018), MCN (Luo et al., 2020), TransVG (Deng et al., 2021), SimREC (Luo et al., 2022), SeqTR (Zhu et al., 2022), LiteREC (Guan et al., 2023a), RefTrans(Li and Sigal, 2021) and our proposed Potamoi are included. For the RES, we introduce MCN, SeqTR, VLT (Ding et al., 2022), RefTrans and our Potamoi for evaluation. Furthermore, to equally evaluate these models on FVG, we add the same radar backbone and the same fusion methods as Potamoi for image and radar to each model. For our Potamoi, we adopt the BERT tokenizer with a maximum token length of 50. Besides, the pretrained weights of FastViT-SA12 (image-backbone) on ImageNet-1K are loaded during training. The head number of proposed MHSCA are 4, 4 and 8 for three stages while the length of sensor agent $n$ is set as 49.
Training. We resize the input image and RVP map as 640 $\times$ 640 (px). We train all models for 80 epochs with a batch size of 16 based on mixed-precision mode. We set the initial learning rate as 1e-3 with the cosine scheduler. We adopt Adam as the optimizer with a momentum of 0.937 and a weight decay of 5e-4. All training experiments are on four RTX A4000 GPUs.
Evaluation metrics. In contrast to conventional REC and RES tasks where a single text prompt refers to a single target, our WaterVG involves a single text prompt referring to multiple targets. Therefore, for the REC task, we choose to jointly evaluate using $AP_{50}$ (Average Precision at IoU 0.5), $AP_{50-95}$ and $AR_{50-95}$ (Average Precision and Recall from IoU 0.5 to 0.95). $AP_{50}$ considers the localization accuracy above 0.5 IoU confidence, while $AP_{50-95}$ is the average accuracy over a range of IoU thresholds from 0.5 to 0.95. It further considers the localization accuracy under different IoU thresholds. $AR_{50-95}$ is the average recall over the IoU threshold from 0.5 to 0.95. It focuses on the coverage of the model to the true target. Meanwhile, for RES tasks, we adopt mIoU to evaluate models. In tables for comparison, bold text indicates the best, and underlined text indicates the second best.
Table 1. Overall comparison of models on WaterVG.
Vision-based Visual Grounding (VVG)
Methods Val Test
@@
MattNet 59.6 31.2 39.7 54.81 57.2 30.2 38.9 54.72
MCN 69.9 39.3 50.7 60.16 66.7 39.4 46.2 59.05
TransVG 70.1 40.7 51.0 - 67.2 39.8 48.0 -
SimREC 56.7 31.3 39.8 - 54.0 30.1 37.9 -
SeqTR 71.2 42.0 51.8 62.98 69.1 41.7 49.2 60.03
LiteREC-S2 65.5 37.9 44.9 - 63.5 36.3 46.3 -
RefTrans 70.8 42.8 52.9 63.72 70.8 42.5 50.1 61.03
VLT - - - 64.13 - - - 61.79
Potamoi 71.6 43.3 54.9 63.16 70.0 42.6 50.7 60.36
Fusion-based Visual Grounding (FVG)
MattNet 63.9 38.1 44.0 57.03 59.1 32.3 38.9 60.32
MCN 70.6 45.6 52.5 61.72 68.1 42.5 46.9 66.23
TransVG 71.8 47.1 54.0 - 69.8 43.6 52.0 -
SimREC 60.5 36.8 42.5 - 56.1 30.9 39.7 -
SeqTR 72.6 47.7 54.2 64.96 70.0 43.5 52.6 68.79
LiteREC-S2 68.9 42.7 47.6 - 64.2 36.3 44.5 -
RefTrans 73.1 47.1 53.8 66.19 69.7 43.1 52.7 70.19
VLT - - - 67.59 - - - 70.90
Potamoi 72.8 47.5 55.0 66.27 70.1 44.8 53.0 70.81

### v2

@@ 3.2. Dataset Statistics
Figure 3 presents the statistics of WaterVG from perspectives of waterway, sentence pattern and query type. The waterway scenario proportion in WaterVG, includes canal, river, lake and moat. Besides, sentence patterns for prompts, including statement (egocentric), question, statement (imperative) and exclamation. Prompt type proportion contains 4 main types of query by exact features, partial features, number and category, respectively. Figure 3 also illustrates the distribution statistics of referred target number, prompt length per sample, referred category distribution and word cloud. In WaterVG, firstly, each text prompt refers to 1-15 targets, with the majority consisting of 1-2 targets. Secondly, The categories of referred target include piers, buoys, sailors, ships, boats, pleasure boats, kayaks and rubbishes, a total of 34,987 referred targets. Thirdly, the distribution of text prompt lengths, ranging from 2 to 40, encompass simple phrases and detailed sentences. Fourthly, the word cloud displays high-frequency words. The above demonstrates the diversity and comprehensiveness of our WaterVG dataset regarding waterway visual grounding, which is primarily reflected in the diversity of waterway scenes, richness of textual prompts in terms of content, patterns, and types, diversity of mentioned target categories, and flexibility in mentioned target number.
@@ 4. Methods
4.1. Task Formulation
Potamoi is a one-stage multi-task visual grounding model for two tasks, REC and RES. Given triple inputs of an image $I\in R^{3\times H_{I}\times H_{W}}$, a frame of radar point clouds $RPC\in R^{5\times N_{R}}$, containing the features $\{x,y,z,v,p\}$ where $v$ and $p$ denote compensated velocity and reflected power, and a text prompt sequence $T=\{w_{1},w_{2},\dotsc,w_{i}\}$ with a length of $N_{T}$. The output of REC is a set of bounding boxes. The $i$-th bounding box is formulated as $B_{i}=\{x_{i},y_{i},w_{i},h_{i},c_{i}\}$, which represents the center coordinates, width, height and IoU confidence of bounding box. RES is to predict a mask of referred targets $M_{RES}\in R^{2\times H_{I}\times H_{w}}$, where the channel dimension contains referred foreground target and unreferred background.
@@ 4.2. Overall Pipeline
Feature Encoders. Firstly, the image encoder adopts FastViT-SA24 (Vasu et al., 2023) with the cascade of reparametrized RepMixer and Multi-Head Self-Attention. Secondly, to prepare for fusion with image, we project 3D radar point clouds from radar to image coordinate, and obtain a 3-channel RVP map, where each position contains the target Range (R), compensated Velocity (V) and reflected Power (P). We adopt the combination of convolution, batch normalization and ReLU activation as the basic block to exact radar features. Each stage of radar encoder contains three blocks, where the convolution kernels are $3\times 3$, $1\times 1$ and $3\times 3$ with the stride of $1$, $1$ and $2$ for local feature extraction, feedforward and downsampling. Whatever image or radar encoder, there are one stem layer and four stages, and the stem layer downsamples the spatial size four times. The feature dimension of stage $i\in\{1,2,3,4\}$ is $(64\times 2^{i-1})\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$.
Feature Fusion. To ensure precise guidance of prompt for both sensors to locate referred targets, the fusion with text requires high-quality integration of features from both sensors. Hence, we adopt a divide-and-conquer strategy called Phased Heterogeneous Modality Fusion. Firstly, due to water-surface reflection and multi-path effect from buildings on the riverbank, radar features unavoidably contain clutter in some scenarios, which results in a certain level of interference in fused sensor features. Secondly, not all prompts contain radar features of targets, so adaptive channel weighting matters. To tackle above problems, a lightweight module called Adaptive Radar Weighting is devised for clutter alleviation, where the last three stage features from the radar encoder will first go through three ARW modules before fusing with image features. Then, three-stage sensor features act as queries to fuse with the keys and values of textual features from the pretrained ALBERT (Lan et al., 2020) by our proposed low-power MHSCA and output three fusion features, which mitigates the quadratic complexity and high power consumption caused by vanilla MHCA (Wu et al., 2023). A lite Path Aggregation Network (PANet) (Li et al., 2022) is to fuse features with various receptive fields, and outputs four-stage pyramid features from $P2$ to $P5$.
@@
Prediction Heads. We feed all four pyramid features to REC heads, which are adapted upon YOLOv8 decouple head (Jocher et al., 2023). Each pyramid feature corresponds to one REC head, containing confidence and bounding box regression branch. The shape of REC output for the stage $i\in\{2,3,4,5\}$ is $(1+4\times RegMax)\times\frac{I_{H}}{2^{i+1}}\times\frac{I_{W}}{2^{i+1}}$, where $1$ is the confidence score and $4$ is four coordinates of bounding box. $RegMax$ is the hyperparameter of Distributed Focal Loss (DFL) (Li et al., 2020). Further, following the paradigm of PP-LiteSeg (Peng et al., 2022), we feed $P2$, $P3$ and $P4$ with SPPM (Jocher et al., 2023) to the RES head, which is based on FLD (Peng et al., 2022). The output shape of RES head is $2\times I_{H}\times I_{W}$, including referred foreground and unreferred background categories.
@@ 4.3. Phased Heterogeneous Modality Fusion
Figure 7. The structure of Slim Cross Attention.
(a) Adaptive Radar Weighting. ARW is employed to dynamically weigh radar features on both channels and space, obtaining high-quality sensor features fused upon vision and radar features while adaptively aligning useful radar features within characteristics of targets contained in the text prompts.
As Figure 6 presents, assuming the radar feature is $F_{R}\in R^{C\times H\times W}$, we exert a linear projection with GeLU activation on it to obtain $\hat{F_{R}}\in R^{C\times H\times W}$. The resulting feature map is then prepared for adaptive channel attention matrix $W_{CA}\in R^{C\times 1\times 1}$ by applying adaptive average pooling to collapse spatial dimensions and computing attention weights by softmax. $\hat{F_{R}}$ is position-wise multiplication with $W_{CA}$ to obtain the channel-weighting feature $F_{R}^{CW}\in R^{C\times H\times W}$. Then spatial weighting matrix $W_{SA}\in R^{1\times H\times W}$ is achieved by convolving $F_{R}^{CW}$ with a GELU activation and a 1 $\times$ 1 convolution, followed by element-wise multiplication with the original features and obtain the final radar feature $F_{R}^{ARW}\in R^{C\times H\times W}$ with both channel and spatial adaptive weighting. Equation 1 and 2 present the whole process above.
@@
(b) Multi-Head Slim Cross Attention. After obtaining three-stage sensor features ($F^{2}_{S}$, $F^{3}_{S}$ and $F^{4}_{S}$) containing contextual features of vision and prior target features captured by radar, there are three linear projection branches from ALBERT to align the channel dimensions of three-stage sensor features. To effectively fuse two heterogeneous modalities with a low cost on USV-based edge devices, we propose Multi-Head Slim Cross Attention (MHSCA) (Figure 7). Given a sensor feature $F_{S}\in R^{C\times H\times W}$ and a text feature $F_{T}\in R^{L\times C}$, we first exert learnable encodings (Dosovitskiy et al., 2020)(Vaswani et al., 2017), $P_{I}$ and $P_{T}$ on them, respectively. Then we adopt 2D depth-wise convolution with $1\times 1$ kernel ($W_{Q}$) to obtain the sensor query ($Q$) while two 1D depth-wise convolutions ($W_{K}$ and $W_{V}$) are for generate key ($K$) and value ($V$). Equation 3 presents the aforementioned process.
(3) $Q=F_{S}W_{Q}+P_{S},K=F_{T}W_{K}+P_{T},V=F_{V}W_{V}+P_{T},$
$Q\in R^{C\times H\times W},K\in R^{L\times C},V\in R^{L\times C},$
Further, we employ adaptive pooling to downsample the sensor query $Q$ (Han et al., 2023), preserving its essential context while reducing its spatial size to obtain the sensor agent $A\in R^{C\times h\times w}$. As the agent of $Q$, $A$ is firstly flattened along spatial dimension as $\hat{A}\in R^{n\times C}$ with $n=h\times w$, and conducts softmax attention $\sigma(\cdot)$ with $K$ and $V$ to calculate the agent-conditional Value matrix $V_{A}$. (Equation 4, 5)
(4) $\hat{A}=Flat(A)^{T},\hat{A}\in R^{n\times C},$
(5) $V_{A}=\sigma(\hat{A},K,V)=\frac{\hat{A}K^{T}}{\sqrt{d}}\cdot V,$
@@
Then, $V_{A}$ is forwarded to the second softmax attention, broadcasting the inter-modal similarity embedded in text feature to the sensor query $Q$, where $A$ acts as the key and $V_{A}$ is the value (Equation 6, 7). Here $Q$ is also firstly flattened as $\hat{Q}\in R^{N\times C}$ with $N=H\times W$.
(6) $\hat{Q}=Flat(Q)^{T},\hat{Q}\in R^{N\times C},$
(7) $F_{TS}=\sigma(\hat{Q},A,V_{A})=\frac{\hat{Q}A^{T}}{\sqrt{d}}\cdot V_{A},$
@@
(8) $\hat{F}_{TS}=(Reshape(F_{TS})+F_{S})W,F_{TS}\in R^{C\times H\times W},$
@@ 4.4. Training Objectives
We reformulate the REC and RES tasks as the bounding box regression and semantic segmentation. Though all targets are denoted as one category, referred target, implicit unbalance between target categories still exist. Hence, as Equation 9 presents, we introduce Distributed Focal Loss (DFL) (Li et al., 2020) to alleviate the unbalanced distribution while CIoU (Zheng et al., 2020) is added to refinedly regress the bounding box size by taking its aspect ratio into account.
@@
For RES task, we combine the dice loss (Sudre et al., 2017) with focal loss (Lin et al., 2017) to jointly subtly optimize the mask prediction (Equation 10).
@@
We empirically argue that the loss functions are not in the same order of magnitude for REC and RES during multi-task training. To avoid the domination of optimization by the task with a large loss, we introduce uncertainty weighting (Kendall et al., 2018) to balance multi-task optimization. Equation 11 demonstrates the loss function during the joint optimization of REC and RES.
@@ 5.1. Settings of Dataset and Implementation
Datasets. Our proposed WaterVG is for training, validation and testing on all models. WaterVG includes two benchmarks, Fusion-based Visual Grounding (FVG) and Vision-based Visual Grounding (VVG). Moreover, to comprehensively valid the performance of Potamoi, we also evaluate our Potamoi on RefCOCO (testA, testB), RefCOCO+ (testA, testB) and RefCOCOg (test-umd) (Yu et al., 2016).
Models. For the REC, MattNet (Yu et al., 2018), MCN (Luo et al., 2020), TransVG (Deng et al., 2021), SimREC (Luo et al., 2022), SeqTR (Zhu et al., 2022), LiteREC (Guan et al., 2023a), RefTrans(Li and Sigal, 2021) and our proposed Potamoi are included. For the RES, we introduce MCN, SeqTR, VLT (Ding et al., 2022), RefTrans and our Potamoi for evaluation. Furthermore, to equally evaluate these models on FVG, we add the same radar encoder and ARW as Potamoi for image and radar to each model. For Potamoi, we adopt ALBERT tokenizer (Lan et al., 2020) with a maximum token length of 50. Besides, the pretrained weights of FastViT-SA24 (image encoder) and other models’ backbones on ImageNet-1K (Deng et al., 2009) are loaded during training. The head number of proposed MHSCA are 4, 4 and 8 for three stages while the length of sensor agent $n$ is set as 49. On RefCOCOs, our Potamoi is trained from scratch.
Training. We resize the input image and RVP map as 640 $\times$ 640 (px). We train all models for 80 epochs on WaterVG with a batch size of 16 while 100 epochs with 64 batch size on RefCOCOs by mixed-precision mode. We set the initial learning rate as 1e-3 with the cosine scheduler. We adopt SGDM as the optimizer with a momentum of 0.937 and a weight decay of 5e-4. All training experiments are on four RTX A4000 GPUs.
Evaluation metrics. Compared with datasets (Yu et al., 2016; Plummer et al., 2015) where a prompt refers to one target, our WaterVG involves a single text prompt referring to multi-targets. Hence, to comprehensively evaluate model performances, we use $AP_{50}$ (Average Precision above IoU 0.5), $AP_{50-95}$ and $AR_{50-95}$ (Average Precision and Recall from IoU 0.5 to 0.95) in REC. For RES tasks, we adopt mIoU to evaluate models. Moreover, we evaluate models on their power consumption (Tu et al., 2023). In tables for comparison, bold text indicates the best, and underlined text indicates the second best. For RefCOCOs, we use $AP_{50}$ as the metric of REC while mIoU as the RES metric.
Figure 8. Images with projected radar point clouds (PCs) (first row) and prediction results by Potamoi (second row) and MCN (Luo et al., 2020) (third row) under different situations. Red bounding boxes are REC predictions while translucent red masks are RES predictions. Blue PCs denote clutter while others represent targets. The gray letters on PCs refers to the distance (m) from targets to the ego USV.
Table 1. Overall comparison of models on WaterVG. MT means multi-task while $P$ is power consumption calculated by hardware API.
Vision-based Visual Grounding (VVG) Fusion-based Visual Grounding (FVG)
Methods MT Val Test Val Test
@@
MattNet 59.6 31.2 39.7 54.81 57.2 30.2 38.9 54.72 - 63.9 38.1 44.0 57.03 59.1 32.3 38.9 60.32 -
MCN 69.9 39.3 50.7 60.16 66.7 39.4 46.2 59.05 81.34 70.6 45.6 52.5 61.72 68.1 42.5 46.9 66.23 77.72
TransVG 70.1 40.7 51.0 - 67.2 39.8 48.0 - - 71.8 47.1 54.0 - 69.8 43.6 52.0 - -
SimREC 56.7 31.3 39.8 - 54.0 30.1 37.9 - 94.42 60.5 36.8 42.5 - 56.1 30.9 39.7 - 88.70
SeqTR 71.2 42.0 51.8 62.98 69.1 41.7 49.2 60.03 90.61 72.6 47.7 54.2 64.96 70.0 43.5 52.6 68.79 84.65
LiteREC 65.5 37.9 44.9 - 63.5 36.3 46.3 - 70.22 68.9 42.7 47.6 - 64.2 36.3 44.5 - 65.79
RefTrans 70.8 42.8 52.9 63.72 70.8 42.5 50.1 61.03 92.97 73.1 47.1 53.8 66.19 69.7 43.1 52.7 70.19 87.21
VLT - - - 64.13 - - - 61.79 - - - - 67.59 - - - 70.90 -
Potamoi 71.6 43.3 54.9 63.16 70.0 42.6 50.7 60.36 66.10 72.8 47.5 55.0 66.27 70.1 44.8 53.0 70.81 61.48
Table 2. Potamoi under waterways, sentence and query patterns.
Waterways Canal Moat River Lake
AP50 72.5 69.4 70.5 67.5
mIoU 72.90 71.81 73.45 66.48
Sentences SE1 SI2 QU3 EX4
AP50 73.6 70.0 65.5 70.1
mIoU 74.12 71.53 66.41 68.74
Queries Exact-Feat Partial-Feat Number Category
AP50 73.1 70.1 62.2 68.7
mIoU 74.35 71.76 63.57 66.87
1. Statement (egocentric), 2. Statement (imperative), 3. Question, 4. Exclamation.